A clock gives one ring to respect 1o'clock, two rings to 2 o'clock and so on.If so, how many times it rings in 24 hours?
A)24
B)300
C)156
D)166

Grade:10

3 Answers

Arun
3 years ago
The clock strikes once at 1 o'clock,twice at 2 o'clock and so on.
[1 day = 24 hours] clock has 12 hours so we have to multiply the series by 2 to make it 24 hours.
 
2 [1 + 2 + 3 + ............ + 12]
 
Sn=n /2 [ a + L ]
Here, a = 1, n = 12, L = 12
 
S12 = 2 x (n/2) [a+L]
S12= 2 x (12/2) [1+12]
S12 = 12 × 13
S12 = 156
 
Hence, the clock will strike 156 times in a day.
 
 
Janani Rajesh Kumar
3 years ago
The answer os the question would be © 156 times
The clock strikes once at 1o’ clock, twice at 2o’ clock and so on. 
So, the number of times the clock will strike in 12 hrs will be:
1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 + 9 + 10 + 11 + 12 = 78 times.
Since the question is asked for 24 hrs you have to multiply this sum by 2 as it continues again from 1 to 12.
Hence the answer would be 78*2 = 156 times.
Vikas TU
3 years ago
Dear student
The clock strikes once at 1 o'clock,twice at 2 o'clock and so on. [1 day = 24 hours] clock has 12 hours so we have to multiply the series by 2 to make it 24 hours. Hence, the clock will strike 156 times in a day